AGW3 + LUCI1 checkout (20180609 UT first half)

Observers: JHill, OKuhn, JPower, JWilliams (Tucson)
Telescope Operator: GBechetti (LBTO)
AO Support: GTaylor (Tucson)
Telescope Support: RHansen (LBTO)
Software Support:
Instrument Support: JMorris (LBTO)
Engineering Support: DCox


We never opened due to clouds and the threat of precipitation. Petr installed the patch which, as he described in an email to telescopework should fix:
  • copointing problems:
    • the copointing limits are now being calculated from simulated M1 tip/tilt values.
  • problems with PCS 'dying' and 'freezing'. (The 'freezing' issue he said happens when the PCS is stopped, so it may not have been encountered in operations).
  • And he noted that NS observations can be carried out with update from ephemeris enabled or disabled (of course must be enabled when the object's rates are changing rapidly) and copointing limits on and off without problems.
We finally issued a closed-dome binocular LUCI preset with fake guide stars to mimic the 1.5" long-slit NS acquisition tried on 20180505 when copointing limits were hit. But there was a MOS error on LUCI1, and that is still unresolved. The connection between the MOS unit and the SW failed during the mask motion. Igor ran some diagnostics, and it seems like it might be a blown fuse. Dave reopened IT #5473, and James and John M will be working to resolve the problem this weekend.

The Plan

AGW3 alignment on 20180603 UT was disrupted because the AGW3 wfs camera could not detect any light. The AGW3 wfs camera was replaced with the spare wfs camera on the afternoon of 20180605. A few nights ago (20180607 UT), in twilight, it was verified that the spare WFS camera on AGw3 does see light. The pupil falls entirely on the lenslet array for on-axis stars, and, for the one off-axis guide star tried, it falls mostly on the lenslet array.

Tonight's plan will the same as that laid out for Com20180603:
  • The AO checkout will be run through at the start
  • Petr is testing a PCS patch to correct copointing limit calculations closed-dome. On-sky tests may follow using the LUCIs on bright sidereal and NS targets.
    • one test will mimic the problem seen 1 month ago, where the sync offset needed to move the target to the 1.5" slit.
  • AGW3 checkout tasks
    • adjust guidecam_hotspot
    • adjust WFScam_hotspot
    • measure on-axis pupil position and update GCS and IDL parameters (WCScam_hotspot_x,_y) Use gcs_wfsc_display and GCSGUI to check.
    • confirm signs for dithering and WFSing
    • measure rotator center (using FS27)
    • measure rotator zeropoints using MOS masks for StoneJ1 and StoneL5P
    • collect transform data


Opening delayed by clouds and virga outflow from thunderstorms in Mexico.

  • Petr installed the PCS patch and did some checks on the simulator.
  • We enabled fake guide stars and sent a closed dome preset to 360 Carlova (1893N), trying to mimic the 1.5" slit acquisition.
    • LUCI1 MOS error in turnout to storage when moving from 0.75 to 1.5/2 slit. Dave T says it looks like a L1 MOS disconnect error.
    • James R, John M and Dave T are all working on it. James says that the telescope can be moved.

07:11: We never opened, so the AGW3 checkout is delayed again. We also have a critical LUCI1 MOS issue (IT 5473 was re-opened).

-- %USERSIG{OlgaKuhn - 2018-06-09}%
Topic revision: r7 - 18 Jul 2018, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback